OpenStudy (anonymous):

Will fan and medal 6x-1 3 _______ = _____ 5 1 @sammixboo

sammixboo (sammixboo):

Well, 3/1 is the same as 3, so let's change that. \(\dfrac{6x-1}{5}=3\)

DivineSolar (divinesolar):

Here i solved it for you, all explained in the png.

sammixboo (sammixboo):

Now we can split up the fraction to make things easier \(\dfrac{6}{5}x~-~\dfrac{1}{5}=3\)

OpenStudy (anonymous):

another?

DivineSolar (divinesolar):

Sammix congratulations on getting anarchy or anarchist cant remember what it is called, oh yes! Ambassador.

OpenStudy (anonymous):

@DivineSolar

DivineSolar (divinesolar):

Yes?

OpenStudy (anonymous):

yay one sec

DivineSolar (divinesolar):

If you are posting another question please post in another section as 1 question per post should be it.